OpenStudy (anonymous):

The roots of 2kx^2+6 = x^2+8x are equal what is the value of k?

OpenStudy (raden):

hint : D = 0 with D = b^2 - 4ac

OpenStudy (anonymous):

why D?

OpenStudy (anonymous):

pls teach me

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

that's the rule if the roots are same then the value of discriminant must be equal 0

OpenStudy (anonymous):

pls write the solution :)

OpenStudy (raden):

first given 2kx^2+6 = x^2+8x subtract on both sides by (x^2+2x), we get 2kx^2+6 - (x^2+8x) = x^2+2x - (x^2+2x) 2kx^2+6 - x^2- 8x = 0 (2k - 1)x^2 - 8x + 6 = 0 looks that the value of a = 2k-1, b = -8, and c = 6 now D = 0 b^2 - 4ac = 0 (-8)^2 - 4 * (2k - 1)(6) = 0 64 - 48k + 24 = 0 solve for k

OpenStudy (anonymous):

thanks :)

OpenStudy (raden):

welcome :)

OpenStudy (anonymous):

k= 11/6? am I right?

OpenStudy (anonymous):

@RadEn am I right? correct answer 11/6?

OpenStudy (raden):

64 - 48k + 24 = 0 88 - 48k = 0 88 = 48k k = 88/48 = 11/6 you are right :D

OpenStudy (anonymous):

THANK YOU VERY MUCH :D

OpenStudy (raden):

my pleasure can help :D